The price elasticity of demand for the services of Kim Jones, MD, is −4.0. The price elasticity of demand for physicians’ services overall is −0.1.

a. Why is demand so much more elastic for the services of Dr. Jones than for the services of physicians in general? (please be detailed)

b. If Dr. Jones reduced prices by 10 percent, how much would volume and revenue change?

c. Suppose that all the physicians in the area reduced prices by 10 percent. How much would the total number of visits and revenue change?

d. Why does it make sense that your answers to questions b and c are so different?

SHOW FORMULAS AND WORK
Answer #1

a)

In this case services of Kim Jones, MD are narrowly defined while services of overall physicians are broadly defined.

In case of services of Kim Jones, MD, it is easier to find close substitutes. At the same time, it is difficult to find a close substitute for a physician's service as a whole. Availability of substitutes makes the demand for Kim Jones, MD relatively elastic. Non availability of close substitutes makes the services of physicians (as a whole) relatively inelastic.

b)

Change in prices=-10%

Change in volume=Price elasticity of demand*Change in prices=-4*(-10%)=40%

Reduction in prices by 10% will lead to an increase of 40% volume in case of Kim Jones, MD.

Change in Revenue=(1+Change in Price)*(1+Change in volume)-1=(1-10%)*(1+40%)-1=26%

Reduction in prices by 10% will lead to an increase of 26% in revenue in case of Kim Jones, MD.

c)

Change in prices=-10%

Change in volume=Price elasticity of demand*Change in prices=-0.1*(-10%)=1%

Reduction in prices by 10% will lead to an increase of 1% volume in case of all physicians.

Change in Revenue=(1+Change in Price)*(1+Change in volume)-1=(1-10%)*(1+1%)-1=-9.1%

Reduction in prices by 10% will lead to a decrease of 9.1% in revenue in case of all physicians.

d)

Following an reduction in prices, there is an increase in volume in both cases. It can be explained with law of demand i,e. keeping the other factors same, quantity demanded increases as price decreases and quantity demanded decreases as price increases.

Total revenue has increased in case of Kim Jones, MD as demand is elastic. Elastic demand means that percent change in price will lead to a higher percent change in volume. 10% reduction in prices leads to 40% increase in volume in this case, causing the total revenue to increase.

Total revenue has decreased in case of all physicians as demand is inelastic. Inelastic demand means that percent change in price will lead to a lower percent change in volume. 10% reduction in prices leads to 1% increase in volume in this case, causing the total revenue to decrease.